OpenStudy (anonymous):

How would I solve Integral of 7^(2x) dx? The indefinite Integral

OpenStudy (gorv):

let u=2x du=2*dx dx=du/2

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (gorv):

\[\int\limits_{}^{}7^u * \frac{ du }{2}\]

OpenStudy (anonymous):

:)